# LLaVA-PruMerge: Adaptive Token Reduction for Efficient Large Multimodal Models

### v4

@@ 3.2 Adaptive Important Token Selection via Outlier Detection
2:Refine $\mathbf{Y}$ to $m$ (adaptive) visual token $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$, in which $m\ll n$.
@@
5:Use the outlier detection algorithm IQR to adaptively select $m$ important visual tokens’ index $\{i_{1},\cdots,i_{m}\}$ based on $\mathbf{a}_{[cls]}$ (see Sec. 3.2).
6:(Optional for PruMerge+) Calculate the outlier ratio $r_{o}=\frac{m}{n}$.
7:(Optional for PruMerge+) Spacial-uniformly sample visual tokens with $r_{o}=\frac{m}{n}$, and get $\{i_{1+m},\cdots,i_{2m}\}$.
8:(Optional for PruMerge+) Update the selected tokens index with $\{i_{1},\cdots,i_{m},1_{1+1},\cdots,i_{2m}\}$.
9:for $p=\{i_{1},\cdots,i_{m}\}$ do
@@
11:     Use $\mathbf{y}_{p}$ as cluster center and run $k$-nearest neighbor algorithm to find $k$ similar tokens, with indices $\{j_{1},\cdots,j_{k}\}_{p}$;
@@
14:(see Sec. 3.3).
15:Output a refined stack of visual token $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$.
3.3 Token Supplement via Similar Key Clustering
Following the selection of more informative visual tokens, we next optimize the utilization of the remaining tokens. While pruned tokens may initially seem extraneous, they hold potential value for the perception capabilities of the LLM backbone. This potential arises particularly in cases where an image contains large object parts that dominate the scene. In such scenarios, overly aggressive pruning could inadvertently diminish the model’s ability to represent the image comprehensively.
To address this, we devise a token merging method aimed at enhancing the representational capacity of the selected unpruned tokens. This method involves the strategic fusion of currently pruned tokens, as depicted in Figure 2. To choose the pruned tokens to merge, we need a way to measure similarity between visual tokens. Here we leverage the self-attention mechanism in Vision Transformers (ViT) to metric such a similarity. Specifically, the final layer’s key vector is utilized as the representation of each image token, and thus we use a dot product similarity metric of keys to calculate which tokens have similar visual information [Bolya et al., 2023]. Specifically, the similarity between visual tokens is computed as
which yields $\mathbf{K}\mathbf{K}^{T}(i,j)$ for tokens $i,j$ in vectorization form for the set of all tokens ${1,2,\cdots,n}$.
3.4 PruMerge+: Bridging the Efficiency-Performance Gap
While PruMergeachieves a remarkable reduction in the number of visual tokens—over tenfold compared to the original setup—the process is not without its drawbacks. Specifically, the compression technique, though efficient, introduces a marginal performance discrepancy between the LLaVA model and its PruMerge-optimized counterpart, LLaVA-PruMerge. To address this, we introduce PruMerge+, a refined version that strikes an optimal balance between token reduction and model performance.
PruMerge+ enhances our original method by maintaining the ability to significantly reduce visual token count—by an average of fourfold—without any resultant performance degradation. This improvement is detailed in Algorithm 1, building upon the token selection strategies outlined in Section 3.2. A new aspect of PruMerge+ lies in its enhanced token selection process. Beyond merely focusing on the previously identified important tokens, PruMerge+ extends its reach to encompass additional visual tokens from areas initially deemed less critical. This is achieved through a spatially uniform sampling of visual tokens, guided by a predetermined ratio informed by the distribution of outlier tokens. This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 3(b), thereby minimizing any performance losses while still achieving substantial token reduction.
@@ 3.5 Discussion
Distinction from Existing Token Reduction Methods. Token reduction methods are especially proposed for inferring vision transformers (ViT) efficiently. Block by block, tokens are gradually reduced in number, correspondingly reducing the computation cost in the internal ViT. However, our approach work is not dedigned for efficient ViT, yet it serves as a plug-and-play module after the final outputs of visual encoder (i.e., ViT). There are two benefits of this design. Firstly, while ViTs typically rely on a single class token to represent the input image, enabling them to maintain performance despite a reduction in intermediate tokens, LMMs usually leverage a large stack of visual tokens. This ensures a comprehensive representation of the visual content, preserving the model’s ability to capture nuanced details. Secondly, considering that the bulk of computational demand within LMMs is attributed to the LLM component rather than the ViT, our approach focuses not on the reduction of tokens but on maximizing the informational content of the pruned visual tokens. This strategy addresses the computational challenges inherent in LMMs without compromising the quality of the visual representation.
@@ 4 Experiments
We first show the empirical performance of our approach when applied to LLaVA-1.5 in Sec 4.1. We then demonstrate the effectiveness of each component in our model in Sec 4.3.
@@ 4.1 Main Results
We evaluate on diverse visual question-answering and reasoning benchmarks including VQAv2 [Goyal et al., 2017], ScienceQA [Lu et al., 2022], TextVQA [Singh et al., 2019], POPE hallucination bench [Li et al., 2023b], MME [Fu et al., 2023], and MMBench [Liu et al., 2023c]. As shown in Table 1, our approach achieves comparable performance with LLaVA-1.5 while performing better than previous works such as BLIP2 [Li et al., 2023a] and InstructBLIP [Dai et al., 2023]. Specifically, in POPE and ScienceQA, our approach even shows better performance than LLaVA-1.5. Note that due to the adaptive nature of PruMerge (see Sec. 3.2), the token numbers for various tasks are different (see 4.3.1), and thus we use the average number on numbers of 6 tasks (i.e., 32) for simplicity.
@@ 4.2 Efficiency Analysis
Backbone (T) Time (ms) Memory (G) Activation (G)
@@
To elucidate the computational efficiency afforded by PruMerge, we utilize the roofline-based LLM-Viewer analysis as developed in [Yuan et al., 2024]. Our investigation is grounded in a theoretical scenario tailored to highlight the impact of PruMerge on processing efficiency within LMMs. Consider a typical scenario where an image of dimensions $336\times 336$ pixels is processed using a CLIP-ViT model, resulting in 576 visual tokens. Accompanying this image is a text prompt, assumed to contain 40 tokens for the sake of this analysis. Through the application of PruMerge, we achieve a dramatic reduction in the number of visual tokens, decreasing the original count by approximately 14.4 times in MME/TextVQA to match the token count of the text prompt ($576/14.4\approx 40$). The implications of this reduction are significant, as demonstrated in Table 2, which outlines the computational cost associated with the LMM prefill process. Notably, PruMerge not only enhances the speed of the LLM prefill process by reducing the required floating-point operations (FLOPs) but also contributes to a reduction in computational memory demands.
@@ 4.3.1 Token Sampling Strategy Analysis
Figure 4: The demo of the different visual tokens sampling strategies.
@@
Sequential sampling: We sample $N$ tokens in the flatted visual tokens. For example, the first 40 tokens are sampled for an apple-to-apple comparison, shown in Figure 4 (c).
Spatial sampling: The sampled $N$ tokens are evenly distributed across the image, as shown in Figure 4 (d-h). We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 6 $\times$ 7 (42 tokens), 7 $\times$ 6 (42 tokens), 5 $\times$ 7 (35 tokens), 7 $\times$ 5 (35 tokens), and 4 $\times$ 4 (16 tokens).
Table 3: Token Sampling Strategy Analysis on Different Tasks. Due to the fact that PruMerge can choose the number of visual tokens based on the input image information density, the number of tokens used in different tasks may differ.
@@
LLaVA-1.5 w/ AITS Vicuna-7B 66.5 54.8 75.7 1221.6
LLaVA-1.5 w/ AITS & TS Vicuna-7B 68.5 56.0 76.3 1350.3
Note that all the experiments are done via the training-free manner. As shown in Table 4, our approach is consistently better than sequential sampling and spatial sampling across all downstream tasks, which demonstrates the effectiveness of the sampling mechanism of LLaVA-PruMerge. Importantly, we observe that LLaVA-PruMerge shows much better performance on TextVQA [Singh et al., 2019]. Such Optical Character Recognition (OCR) task requires detailed information about the text, which demonstrates that LLaVA-PruMerge extracts the key information in the images with enough details. This quantitative result aligns with the visualization of LLaVA-PruMerge attentive tokens in Figure 1(a) (b), where more attentive tokens are distributed on the foreground text in the images.
@@ 4.3.2 Effectiveness of Each Module in PruMerge
LLaVA-PruMerge can be conducted both in training-free and training-needed manners. With fine-tuning, the large language model can adapt to the new structure of visual tokens, which could further enhance the performance on vision-language tasks. As shown in Table 5, with fine-tuning, our approach does bring better performance for diverse tasks, including ScienceQA [Lu et al., 2022], TextVQA [Singh et al., 2019], POPE [Li et al., 2023b], and MME [Fu et al., 2023].
Table 5: Ablation on training free and fine-tuning for our approach. With further fine-tuning, the performance of LLaVA-PruMerge can be further enhanced.
Method LLM SQAI VQAT POPE MME
LLaVA-1.5 Vicuna-7B 66.8 58.2 85.9 1510.7
LLaVA-PruMerge w.o. fine-tuning Vicuna-7B 68.0 54.0 76.2 1250.1
LLaVA-PruMerge w. fine-tuning Vicuna-7B 68.5 56.0 76.3 1350.3
@@ 5 Conclusion
In this paper, we improve the efficiency of Large Multimodal Models (LMMs) from the perspective of reducing the quantity of visual tokens. By leveraging the spatial redundancy in visual tokens, we proposed a plug-and-play token reduction module that employs the similarity between the class token and spatial tokens as a key criterion for pruning and merging visual tokens.
Our approach, applied to LLaVA-1.5, demonstrated that by utilizing only 6.9% of visual tokens on average, the pruned tokens can maintain comparable performance across a wide range of visual question-answering and reasoning tasks. Notably, our work highlights the potential for significant computational savings without sacrificing the reasoning capabilities of LMMs. We hope our work inspires further exploration into the interplay between efficiency and performance in LMMs.
Acknowledgement
This work was supported in part by NSF CAREER IIS2150012, and Institute of Information & communications Technology Planning & Evaluation(IITP) grants funded by the Korea government(MSIT) (No. 2022-0-00871, Development of AI Autonomy and Knowledge Enhancement for AI Agent Collaboration) and (No. RS2022-00187238, Development of Large Korean Language Model Technology for Efficient Pre-training), and Microsoft Accelerate Foundation Models Research Program.

### v5

@@ 3.4 PruMerge+: Bridging the Efficiency-Performance Gap
2:Refine $\mathbf{Y}$ to $m$ (adaptive) visual tokens $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$, in which $m\ll n$.
@@
5:Use the outlier detection algorithm IQR to adaptively select $m$ important visual tokens’ indices $\{i_{1},\cdots,i_{m}\}$ based on $\mathbf{a}_{[cls]}$ (see Sec. 3.2).
6:(Optional for PruMerge+, see Sec. 3.4) Calculate the outlier ratio $r_{o}=\frac{m}{n}$.
7:(Optional for PruMerge+) Spatial-uniformly sample visual tokens with $r_{o}$, and get $\{i_{1+m},\cdots,i_{2m}\}$.
8:(Optional for PruMerge+) Update the selected tokens’ index with $\{i_{1},\cdots,i_{m},i_{m+1},\cdots,i_{2m}\}$.
9:for $p=\{i_{1},\cdots,i_{m}\}$ do (see Sec. 3.3)
@@
11:     Use $\mathbf{y}_{p}$ as cluster center and find the $k$ most similar tokens, with indices $\{j_{1},\cdots,j_{k}\}_{p}$;
@@
14:Output a refined stack of visual tokens $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$.
@@ 3.5 Discussion
Distinction from Existing Token Reduction Methods. Token reduction methods [Liu et al., 2022, Yin et al., 2022, Liang et al., 2022] have been proposed for accelerating ViT computation speed. Block by block, tokens are gradually reduced in number, correspondingly reducing the computation cost in the internal ViT. By contrast, our approach is not specifically designed to increase the efficiency of a ViT. It instead aims to improve the overall efficiency of a large multimodal model, where ViT is only one relatively light cost component as shown in Fig. 1. There are two benefits of this design. First, while ViTs typically rely on a single class token to represent the input image, which enables them to maintain performance despite a reduction in intermediate tokens, LMMs usually require a large stack of visual tokens. This ensures a comprehensive representation of the visual content, preserving the model’s ability to capture nuanced details. Thus, using previous token merging methods to obtain one refined class token as representation of visual input is not consistent with the literature of large multimodal models. Second, considering that the bulk of computational demand within LMMs is attributed to the LLM component rather than the ViT, our approach focuses not only on the reduction of tokens but also on maximizing the informational content of the pruned visual tokens. This strategy addresses the computational challenges inherent in LMMs with minimal compromise in the quality of the visual representation.
@@ 4 Experiments
We first show the empirical performance of our approach when applied to LLaVA-1.5 in Sec 4.1. We then analyze the efficiency improvement by using our PruMerge on LMM in Sec 4.2. To show the generalization ablity, we conduct a series of experiments in Sec. 4.3. Finally, we demonstrate the effectiveness of each component in our model in Sec 4.4.
@@ 4.1 Main Results
We evaluate on diverse visual question-answering and reasoning benchmarks including VQAv2 [Goyal et al., 2017], ScienceQA [Lu et al., 2022], TextVQA [Singh et al., 2019], POPE hallucination bench [Li et al., 2023b], MME [Fu et al., 2023], and MMBench [Liu et al., 2023c]. As shown in Table 1, our approach achieves comparable performance with LLaVA-1.5 despite using only a small fraction of the visual tokens, and performing better than previous works such as BLIP2 [Li et al., 2023a] and InstructBLIP [Dai et al., 2023]. Specifically, in POPE and ScienceQA, our approach even shows better performance than LLaVA-1.5. Note that due to the adaptive nature of PruMerge (see Sec. 3.2), the token numbers for various tasks are different (see 4.4.1), and thus we use the average number on numbers of 6 tasks (i.e., 32) for simplicity.
@@ 4.2 Efficiency Analysis
Backbone (TB) Time (ms) Memory (GB) Activation (GB)
@@
To elucidate the computational efficiency afforded by PruMerge, we utilize the roofline-based LLM-Viewer analysis developed in [Yuan et al., 2024]. Our investigation is grounded in a theoretical scenario tailored to highlight the impact of PruMerge on processing efficiency within LMMs. Consider a typical scenario where an image of dimensions $336\times 336$ pixels is processed using a CLIP-ViT model, resulting in 576 visual tokens. Accompanying this image is a text prompt, assumed to contain 40 tokens for the sake of this analysis. Through the application of PruMerge, we achieve a dramatic reduction in the number of visual tokens, decreasing the original count by approximately 14.4 times in MME/TextVQA to match the token count of the text prompt ($576/14.4\approx 40$). The implications of this reduction are significant, as demonstrated in Table 2, which outlines the computational cost associated with the LMM prefill process. Notably, PruMerge not only enhances the speed of the LLM prefill process by reducing the required floating-point operations (FLOPs) but also contributes to a reduction in computational memory demands.
@@
4.3 Generalization on Video-LLM
To assess the generalization capabilities of PruMerge and PruMerge+ across different modalities, we next extend our approach to Video-LLaVA [Lin et al., 2023]. Video-LLaVA is one of the most popular open-soruced Video-LLMs. We seamlessly integrate both algorithms into Video-LLaVA without the need for additional training, enabling us to bypass re-training on video datasets during inference. The outcome of this integration is shown in Table 3. Video-LLaVA samples 8 frames from a video clip and extracts $8\times 16\times 16=2048$ visual tokens using a visual encoder for LLM perception, which is 4 times of visual token than LLaAV-1.5 [Liu et al., 2023a]. Our Algorthms PruMerge and PruMerge+ can adaptively select important 256 (12.5% on average) and 256 (25.0% on average) important visual tokens, respectively. The results demonstrate that our algorithms not only reduce the number of visual tokens in Video-LLaVA but also is able to enhance its performance. This finding is noteworthy as it suggests a significant redundancy in the visual tokens used by video-LLMs. Exploring ways to further capitalize on this redundancy could shape future research directions.
Table 3: A comparison of different LVMs on video reasoning benchmarks. Like Video-LLaVA, ChatGPT-Assistant is employed to evaluate performance. The version of ChatGPT is “gpt-3.5-turbo”. Note that we directly add PruMerge and PruMerge+ to Video-LLaVA during inference (without training our own model).
Methods LLM size MSVD-QA MSRVT-QA ActivityNet-QA
Accuracy Score Accuracy Score Accuracy Score
FrozenBiLM 1B 32.2 - 16.8 - 24.7 -
VideoChat 7B 56.3 2.8 45.0 2.5 - 2.2
LLaMA-Adapter 7B 54.9 3.1 43.8 2.7 34.2 2.7
Video-LLaMA 7B 51.6 2.5 29.6 1.8 12.4 1.1
Video-ChatGPT 7B 64.9 3.3 49.3 2.8 35.2 2.7
Video-LLaVA 7B 70.7 3.9 59.2 3.5 45.3 3.3
Video-LLaVA + PruMerge 7B 71.1 3.9 58.4 3.5 48.3 3.4
Video-LLaVA + PruMerge+ 7B 71.1 3.9 59.3 3.6 47.7 3.4
@@ 4.4.1 Token Sampling Strategy Analysis
Sequential sampling: We sample $N$ tokens in the flatted visual tokens. For example, the first 40 tokens are sampled for an apples-to-apples comparison, shown in Figure 4 (c).
Figure 4: Different token sampling strategies.
Spatial sampling: The sampled $N$ tokens are evenly distributed across the image, as shown in Figure 4 (d-h). We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 5 $\times$ 7 (35 tokens), 7 $\times$ 5 (35 tokens), and 4 $\times$ 4 (16 tokens).
Table 4: Token Sampling Strategy Analysis on Different Tasks.
@@
LLaVA-1.5 w. AITS Vicuna-7B 66.5 54.8 75.7 1221.6
LLaVA-1.5 w. AITS & TS Vicuna-7B 68.5 56.0 76.3 1350.3
Table 6: Ablation on training free and fine-tuning for our approach. With fine-tuning, the performance of LLaVA-PruMerge can be further enhanced.
Method LLM SQAI VQAT POPE MME
LLaVA-1.5 Vicuna-7B 66.8 58.2 85.9 1510.7
LLaVA-PruMerge w.o. LoRA-FT Vicuna-7B 68.0 54.0 76.2 1250.1
LLaVA-PruMerge w. LoRA-FT Vicuna-7B 68.5 56.0 76.3 1350.3
Note that all the experiments are done via a training-free manner. As shown in Table 6, our approach is consistently better than sequential sampling and spatial sampling across all downstream tasks, which demonstrates the effectiveness of the sampling mechanism of LLaVA-PruMerge. Importantly, we observe that LLaVA-PruMerge shows much better performance on TextVQA [Singh et al., 2019]. Such Optical Character Recognition (OCR) task requires detailed information about the text, which demonstrates that LLaVA-PruMerge extracts the key information in the images with enough details. This quantitative result aligns with the visualization of LLaVA-PruMerge attentive tokens in Figure 1(b), where more attentive tokens are distributed on the foreground text in the images.
@@ 4.4.2 Effectiveness of Each Module in PruMerge
Finally, LLaVA-PruMerge can be conducted in either a training-free or fine-tuning manner. With fine-tuning, the large language model can adapt to the new structure of visual tokens, which could further enhance the performance on vision-language tasks. As shown in Table 6, with fine-tuning, our approach does bring better performance for diverse tasks, including ScienceQA [Lu et al., 2022], TextVQA [Singh et al., 2019], POPE [Li et al., 2023b], and MME [Fu et al., 2023].
@@ 5 Conclusion
In this paper, we improve the efficiency of Large Multimodal Models (LMMs) from the perspective of reducing the quantity of visual tokens. By leveraging the spatial redundancy in visual tokens, we proposed a plug-and-play token reduction module that employs the similarity between the class token and spatial tokens as a key criterion for pruning and merging visual tokens. Our approach, applied to LLaVA-1.5, demonstrated that by utilizing only 6.9% of visual tokens on average, the pruned tokens can maintain comparable performance across a wide range of visual question-answering and reasoning tasks. Notably, our work highlights the potential for significant computational savings without sacrificing the reasoning capabilities of LMMs. We hope our work inspires further exploration into the interplay between efficiency and performance in LMMs.
Appendix A Appendix
A.1 Limitations and Future Work
Our exploration of LLaVA-PruMerge has two primary limitations. First, the visual token compression, while efficient, is not entirely lossless. This results in a marginal performance gap between the original model (LLaVA) and our optimized version LLaVA-PruMerge. We are committed to advancing our research towards achieving a fully lossless token compression algorithm, aiming to close these performance gaps. Second, the scope of our validation efforts is somewhat constrained by the computational resources typically available in academic settings. This limitation has precluded a comprehensive assessment of PruMerge’s applicability to larger-scale models, such as those envisioned in the LLaVA-Next [Liu et al., 2024] framework with a Yi-34B backbone. Future investigations will seek to extend our methodology to these more expansive model architectures, exploring its potential for generalization and broader impact.
A.2 Societal Impacts
In this study, we propose a technique that improves the efficiency of Large Multimodal Models (LMMs), making them more accessible. This approach helps to democratize LMMs by lowering deployment costs and hardware barriers, facilitating their use in edge computing. However, it does not mitigate the potential misuse of LMMs by malicious actors.